What is the sum of squares for the following sample of 5 scores: 5, 4, 3, 9, 4?
A. 5
B. 15
C. 20
D. 22
Answer #1

From given 5 scores we get, sample mean is,

5+4+3+9+4 25

We want to find, the sum of squares (z -T)

(z-z)2 = (5-5)2 + (4-5)2 + (3-5)2 + (9-5)2 + (4-5)2

sum ( x -ar x)^2 =0 + ( -1)^2 +( -2)^2 +( 4)^2 +( -1)^2

(z-i)2-0+1+1+1641

sum ( x -ar x)^2 = 22

The sum of squares is 22

Answer: D) 22